# Regimes and Transitions of the Nonlinear Temporal Talbot Effect: Underlying Mechanism for A-Type Breathers, Soliton Crystals, and Soliton Gas

### v1

@@
Regimes and Transitions of the Nonlinear Temporal Talbot Effect: Underlying Mechanism for A-Type Breathers, Soliton Crystals, and Soliton Gas
@@
journal: opticajournalarticletype: Research Article
{abstract*}
A frequency comb from a phase-modulated continuous-wave laser is subject to the temporal Talbot effect and modulational instability in an optical fiber or semiconductor. Using Soliton Radiation Beat Analysis and a dispersion relation for comb lines, I identify distinct regimes of the nonlinear Talbot effect of A-type breathers, soliton crystals, and soliton gas that arise from varying hierarchies of self-phase modulation, cross-phase modulation, and the type of four-wave mixing (regular vs. cascaded). The temporal Talbot effect governs the spatio-temporal distribution of linear and nonlinear pulses and is closely related to the Fermi-Pasta-Ulam-Tsingou recurrence phenomenon. The study advances the fields of Nonlinear Optics and Wave Theory.
@@ 1 Introduction
The frequency comb from a phase-modulated CW laser is subject to two fundamental effects in nonlinear dispersive media such as optical fibers and semiconductors. The first one is the Talbot effect, and the second one is modulational instability (MI). The Talbot effect refers to the self-imaging of a periodic optical pattern under the influence of diffraction or dispersion, both of which are linear effects in the optical field amplitudes [7]. MI, albeit depending on material dispersion, particularly anomalous dispersion, is nonlinear and associated with the dynamical growth and evolution of periodic perturbations on a CW background [8]. Despite their simultaneous action on a frequency comb, these effects have been treated mostly separately; their combined action remains largely ununderstood.
@@
The propagation of a periodic pattern in a medium becomes quickly nonlinear due to light-matter interaction. Being subject to material nonlinearity, the initially linear Talbot effect becomes nonlinear. Studies and utilization of the nonlinear Talbot effect have attracted researchers’ interest for the last two decades. Thus, nonlinear spatial Talbot effect was observed in periodically poled $\text{LiTaO}_{3}$ and $\text{LiNbO}_{3}$ crystals [31], [32], [33], [34] as well as nonlinear photonic crystals [35] opening a pathway for high-resolution lithography, reported as a means for excitation of nonlinear beams such as Akhmediev breathers (AB) and solitons [36], [37], as an origin of the optical-pattern formation in cold atomic clouds [38], and as a way to generate excitation gratings in the hard X-rays [39]. Despite the growing interest in and understanding of the nonlinear Talbot effect, it still requires further study to be effectively utilized in applications. Particularly, the nonlinear temporal Talbot effect lacks systematic studies and thorough understanding
@@
MI has been extensively studied in systems governed by the Nonlinear Schrödinger Equation (NLS) since the 1960s [40]. Those are, for instance, optical fibers, water waves, and the Bose-Einstein condensate. MI’s mathematical treatment has led to a variety of NLS solutions, which have been followed by experimental observations. Thus, the fundamental MI-related solution is the AB, a temporally periodic, spatially localized, nonlinear wave, reported in 1986 [41] and [42]. Other solutions are spatio-temporarily localized Peregrine solitons [43], [44], [45], [46] and their closely related brothers called rogue waves [47], [48]. The recognition of MI’s close relationship to the Fermi–Pasta–Ulam–Tsingou (FPUT) recurrence led to finding AB recurrence under loss as a perturbation [49], construction of spatio-temporally periodic solutions under higher-order MI [40], as well as a quite recent finding and observation of a doubly periodic umbrella solution class to ABs called A-type and B-type breathers [50], [51], [52], [53], [54].
@@
Although a frequency comb from a phase-modulated CW laser is subject to both MI [8] and temporal Talbot effect in fibers [19], the studies connecting these two fundamental effects are scarce. Thus, the nonlinear temporal Talbot effect has been brought into connection with emergence of rogue waves from doubly periodic initial conditions [55], [56], [57] and frequency-comb generation via cross-phase modulation (XPM) [58], [59], both MI-related effects. Further studies are needed to understand better the relationship between the Talbot effect and MI, particularly concerning the spatio-temporal recurrence of optical patterns in fibers and semiconductors, to create an exhaustive theory of the nonlinear temporal Talbot effect.
@@
In the Letter  [60], my colleague Michael Böhm and I reported a strong relationship between the temporal Talbot effect and MI-induced nonlinear-wave generation, studying a phase-modulated CW laser field in optical fibers. We recognized this relationship using Soliton Radiation Beat Analysis (SRBA). This numerical technique allows retrieving soliton content of nonlinear waves generated from arbitrary inputs in optical fibers [61], [62], [63], [64]. Thus, we reported input-power-dependent transitions from a quasi-linear regime, governed primarily by linear Talbot effect at low input powers, to the regimes of the nonlinear Talbot effect of soliton crystals and separated solitons (cf. Fig. 2, left). These regime transitions happened at well-defined threshold input powers, giving rise to the interpretation that the NLS changes the type of its solutions at these values.
@@
To the best of my knowledge, those were the first reported results that connected the temporal Talbot effect with MI-related generation of soliton crystals and solitons in such detail. However, the results were derived rather heuristically from our SRBA observations. Many questions remained open, in particular, a theoretically backed-up explanation of the observed regimes and their transitions.
@@
Here, I build upon the results of Ref. [60] and study the transitions of the temporal Talbot effect and the physical mechanisms behind them using SRBA and bridging SRBA with a derived dispersion relation for frequency-comb lines. It allows me to reveal four input-power regimes: the regime of the linear temporal Talbot effect and the nonlinear regimes of A-type breathers, soliton crystals, and separated solitons (also called soliton gas due to a weak interaction between the solitons, [65]). I show that the (higher-order) interference of the discrete spatial modes of the frequency-comb lines originates from the temporal Talbot effect and governs the spatio-temporal distribution of the optical field in all regimes, allowing for the development of an exhaustive theory of the nonlinear temporal Talbot effect. The present study confirms the observation that frequency-comb lines can encode line-specific solitons, first reported in Ref. [60]. Additionally, I draw a difference between A-type breathers and soliton crystals. Thus, A-type breathers originate from regular FWM at quite low input powers, whereas soliton crystals stem from cascaded FWM at higher input powers. Exhibiting these characteristics, soliton crystals in single-pass optical fibers are no different from soliton crystals in cavities. Last but not least, I show that phase modulation depth is an important parameter that controls the type of nonlinear waves. Thus, a small modulation depth results in ABs under FPUT close to the Peregrine-soliton limit, whereas higher values of modulation depth lead to the formation of doubly periodic solutions such as A-type breathers and soliton crystals.
@@
To conclude, for decades, MI-induced nonlinear wave generation has been considered as a separate field of studies, giving rise to different types of NLS solutions. These solutions are important for the description and classification of various nonlinear waves in optical fibers and semiconductors. However, they provide only a limited understanding of the physics behind. My present study shows that a deeper and more sophisticated insight into the dynamics of a phase-modulated CW laser field stems from the perspective of the nonlinear temporal Talbot effect that integrates SPM and MI-related effects of FWM and XPM in the theory of the well-known linear temporal Talbot effect. With that, it contribute to the advancement of Nonlinear Optics and Wave Theory.
@@
The study is structured as follows. In Sec. 2, I present the system under study, a phase-modulated CW laser field in a single-mode optical fiber, and show that the operation happens beyond the standard AB theory. In Sec.3, I discuss the methodology for theoretical and numerical studies, including the SRBA. In the Results and Discussion section (Sec. 4), I present a dispersion relation for optical-frequency comb lines that includes group-velocity dispersion (GVD) and, with it, the footprint of the linear temporal Talbot effect, SPM, XPM, and FWM, (Sec. 4.1); then, I bridge this dispersion relation with SRBA (Sec. 4.2); I discuss the transitions of the temporal Talbot effect from the linear to nonlinear regimes and draw the difference between the A-type breathers and soliton crystals (Sec. 4.3); I provide a confirmation that frequency-comb lines can encode line-specific solitons ([60], Sec. 4.4); and, finally, I discuss the impact of the modulation depth on the nonlinear-wave evolution (Sec. 4.5). I conclude and present a summary of the achieved results in Sec. 5.
@@ 2.1 The System Under Study: A Phase-Modulated Continuous-Wave Optical Field
The optical system I study here is depicted in Fig. 1 (right). It consists of a CW laser with emission at a central wavelength of $\lambda_{0}=1554.6~\text{nm}.$ The laser output is phase-modulated by the phase modulator (PM) with modulation frequency $\Omega=15.625~\text{GHz}$ and modulation depth $m.$ The phase-modulated CW field is propagated through a single-mode fiber (SMF) and recorded by an optical sampling oscilloscope (OSA) and optical spectrum analyzer (OSA).
@@
Figure 2: Left: Soliton Radiation Beat Analysis of input-power dependent regimes in dB (cf. [60]). Green arrows mark ridges of spatio-temporarily separated solitons. The yellow arrow marks the fundamental Talbot frequency $Z_{T}$ (Eq. 13). Right: Corresponding optical peak power $P(z)=|A(z,t=0)|^{2}$ in W. The PM modulation depth is $\text{m}=1.$ Dashed lines mark input-power dependent transitions of the nonlinear temporal Talbot effect.
According to Ref. [60], these regimes arise due to the temporal Talbot effect, a self-imaging effect in dispersive media such as optical fibers or semiconductors (cf. [26], [27]). The condition for this effect is satisfied by the initial frequency comb that is provided through the phase modulation of a CW laser field [19], [20], [21] (Fig. 1, right). The fiber dispersion acts as a unitary phase filter, inducing the break-up of the phase-modulated CW field into a train of identical pulses [23]. These pulses are subject to the Kerr nonlinearity of the fiber glass, which leads to the formation of the nonlinear regimes of soliton crystals and separated solitons. In Sects. 4.1, 4.2, 4.3, I provide a detailed explanation of how these regimes arise, building a bridge between the linear temporal Talbot effect and the nonlinear SPM and MI-driven FWM and XPM of frequency-comb lines. It allows for a concise theory of the nonlinear temporal Talbot effect.
@@ 2.2 Operation Beyond the Established Akhmediev Breather Theory
Fig. 4 shows the MI gain $g(\Omega)$ and the corresponding AB parameters $a$ and $b$ for input powers studied in Ref. [60] in more detail below. The AB parameters are fundamental to the AB solution of the NLS and are defined as $2a=(1-(\Omega/\Omega_{c})^{2})$ and $b=\sqrt{8a(1-2a)}$ with $\Omega_{c}=4\gamma P_{0}/|\beta_{2}|$ such that the maximum gain happens for $b=1$ and $a=1/4$ [8]. For $\beta_{2}=-23~\text{ps}^{2}/\text{km}$ used here, the AB parameters are $\leq 0$ and, thus, not satisfactory for an AB solution for the most significant input power values of $P_{0}=0.046~\text{W},$ $0.15~\text{W},$ and $0.27~\text{W}$ (cf. Fig. 3, top). Also, the higher-order MI does not come into question as a driving mechanism for our observations, as it requires a value of $a\geq 0.375$ [40] which is achievable only for low absolute values of GVD. Apart from that, the PM modulation frequency RF of $\Omega=15.625~\text{GHz}$ generates only a little MI gain far beyond its input-power dependent maxima. From these observations, I conclude that the results reported in Ref. [60] and studied here in more detail do not originate from ABs under FPUT recurrence despite their visual similarity ([3]). Rather, they are the result of the nonlinear temporal Talbot effect that incorporates SPM and MI-driven FWM and XPM of frequency-comb lines (Sec. 4).
@@
A more generalized class of NLS solutions constitutes the A-type and B-type breathers ([51], [52], [53]). In Ref. [54], the authors discussed that A-type breathers exist beyond the standard MI bandwidth of modulation frequencies, i.e., for higher modulation frequencies than the cut-off frequency of $g(\Omega)$ (Fig. 4, left), which gives rise to "extraordinary" MI. Below, I will discuss a regime of the nonlinear temporal Talbot effect for input powers of $0.046~\text{W}<P_{0}\leq 0.15~\text{W}$ that hosts doubly-periodic waves with properties of A-type breathers (Sec. 4.3). Apparently, the PM modulation frequency does not need to exceed the cut-off of $g(\Omega)$. It is also possible to generate A-type breathers within the MI bandwidth for a small value of $g(\Omega=15.625~\text{GHz}).$
@@
Soliton crystals, spatio-temporally periodic soliton compounds for $0.15~\text{W}<P_{0}\leq 0.27~\text{W},$ constitute a separate class of NLS solutions. The difference to A-type breathers lies in the underlying physics. As discussed in Secs. 4.1, 4.2, 4.5, the A-type breathers are generated by MI-driven regular FWM at quite low input powers and small PM modulation depths; soliton crystals originate from cascaded FWM at higher input powers and strong PM modulation depths. For higher input powers, $P>0.27~\text{W},$ soliton crystals dissolve into spatio-temporarily separated solitons, i.e., soliton gas ([65]).
@@
I will show that higher-order interference between the frequency-comb lines driven by the temporal Talbot effect is the underlying mechanism for any type of the discussed NLS solutions (Sec. 4.3). Therefore, I conclude that an exhaustive consideration of nonlinear-wave evolution in optical fibers should be conducted from the perspective of the nonlinear temporal Talbot effect, rather than considering SPM or MI-driven effects separately.

### v2

# Regimes and Transitions of the Nonlinear Temporal Talbot Effect: Underlying Mechanisms of A-Type Breathers, Soliton Crystals, and Soliton Gas

@@
Abstract
A frequency comb generated from a phase-modulated continuous-wave laser is simultaneously subject to the temporal Talbot effect and modulational instability (MI) when propagating through a piece of optical fiber. The temporal Talbot effect refers to the dispersion-driven self-imaging of optical pulses and is, per se, linear in optical field amplitudes. MI is a nonlinear effect. Despite growing interest and a variety of possible applications, a concise theory of the nonlinear temporal Talbot effect that incorporates nonlinear effects is not yet available; the self-imaging of optical patterns under the influence of nonlinearity remains largely unexplored. Here, I derive a dispersion relation for frequency-comb spatial modes. It integrates the contributions of the linear temporal Talbot effect, self-phase modulation, and MI-driven cross-phase modulation and four-wave mixing (FWM) between frequency-comb lines, paving the way to the development of a concise theory of the nonlinear temporal Talbot effect. With the help of Soliton Radiation Beat Analysis, a technique for retrieving the soliton content of optical fields in fibers, I demonstrate that the temporal Talbot effect seeds the spatio-temporal distribution of the field by driving higher-order spatial-mode interference. Nonlinear effects modify this interference, linking Talbot effect to Fermi–Pasta–Ulam–Tsingou recurrence of optical pulses and producing input-power-dependent regime transitions between the linear Talbot effect, A-type breathers, soliton crystals, and soliton gas. I show that A-type breathers arise from regular FWM at lower input powers, whereas soliton crystals emerge from cascaded FWM at higher powers. This study advances the fields of Nonlinear Optics and Wave Theory.
keywords:
Temporal Talbot Effect , Frequency Combs , Soliton Crystals , Breathers , Optical Solitons , Fermi–Pasta–Ulam–Tsingou (FPUT) Recurrence
journal: Optics & Laser Technology
@@ Abstract
[1]organization=Multitel Innovation Centre,addressline=Rue Pierre et Marie Curie 2, city=Mons, postcode=7000, country=Belgium
{highlights}
Unified theory of the nonlinear temporal Talbot effect is established.
Four input-power regimes identified and explained: linear Talbot, A-type breathers, crystals, gas.
Dispersion relation derived for frequency-comb spatial modes under nonlinearity.
Higher-order Talbot interference shown as seed of nonlinear waves and their recurrence.
Distinction drawn between regular-FWM breathers and cascaded-FWM soliton crystals.
@@ 1 Introduction
The frequency comb from a phase-modulated CW laser is subject to two fundamental effects in nonlinear dispersive media such as optical fibers and semiconductors. The first one is the Talbot effect, and the second one is modulational instability (MI). The Talbot effect refers to the self-imaging of a periodic optical pattern under the influence of diffraction or dispersion, both of which are linear effects in the optical field amplitudes [7]. MI, albeit depending on material dispersion, particularly anomalous dispersion, is nonlinear and associated with the dynamical growth and evolution of periodic perturbations on a CW background [8]. Despite their simultaneous action on a frequency comb, these effects have been treated mostly separately. The combined action of these two effects is not yet fully understood, and a concise theory of the nonlinear Talbot effect is still lacking, one that would integrate nonlinearities into the concepts of the initially linear temporal and spatial Talbot effects.
@@
The propagation of a periodic pattern in a medium becomes quickly nonlinear due to light-matter interaction. Being subject to material nonlinearity, the initially linear Talbot effect becomes nonlinear. Studies and utilization of the nonlinear Talbot effect have attracted researchers’ interest for the last two decades. Thus, nonlinear spatial Talbot effect was observed in periodically poled $\text{LiTaO}_{3}$ and $\text{LiNbO}_{3}$ crystals [31], [32], [33], [34] as well as nonlinear photonic crystals [35] opening a pathway for high-resolution lithography, reported as a means for excitation of nonlinear beams such as Akhmediev breathers (AB) and solitons [36], [37], as an origin of the optical-pattern formation in cold atomic clouds [38], and as a way to generate excitation gratings in the hard X-rays [39]. Despite the growing interest in the nonlinear Talbot effect, spatial and temporal, it still requires further studies to be fully understood and effectively utilized in applications. Whereas some approaches have been undertaken to create a theory of the nonlinear spatial Talbot effect (cf. [36], [37]), a concise theory of the nonlinear temporal Talbot is still missing, to the best of my knowledge.
@@
MI has been extensively studied in systems governed by the Nonlinear Schrödinger Equation (NLS) since the 1960s [40]. Those are, for instance, optical fibers, water waves, and the Bose-Einstein condensate. MI’s mathematical treatment has led to a variety of NLS solutions, which have been followed by experimental observations. Thus, the fundamental MI-related solution is the Akhmediev breather (AB), a temporally periodic, spatially localized, nonlinear wave, reported in 1986 [41], [42]. Other solutions are spatio-temporally localized Peregrine solitons [43], [44], [45], [46] and their closely related brothers called rogue waves [47], [48]. The recognition of MI’s close relationship to the Fermi–Pasta–Ulam–Tsingou (FPUT) recurrence led to finding AB recurrence under loss as a perturbation [49], construction of spatio-temporally periodic solutions under higher-order MI [40], as well as a quite recent finding and observation of a doubly periodic umbrella solution class to ABs called A-type and B-type breathers [50], [51], [52], [53], [54].
@@
Although a frequency comb from a phase-modulated CW laser is Although a frequency comb from a phase-modulated CW laser is simultaneously subject to both MI [8] and temporal Talbot effect in fibers [19], the studies connecting these two fundamental effects to a unified framework are scarce. Thus, the nonlinear temporal Talbot effect has been brought into connection with the emergence of rogue waves from doubly periodic initial conditions [55], [56], [57], and frequency-comb generation via cross-phase modulation (XPM) [58], [59], both MI-related effects. Further studies are needed to better understand the relationship between the temporal Talbot effect and MI and their simultaneous action, particularly concerning the spatio-temporal recurrence of optical patterns in fibers and semiconductors, to create a concise theory of the nonlinear temporal Talbot effect.
@@
In the Letter  [60], my colleague Michael Böhm and I reported a strong relationship between the temporal Talbot effect and MI-related nonlinear-wave generation, studying a phase-modulated CW laser field in optical fibers. We recognized this relationship using Soliton Radiation Beat Analysis (SRBA). This numerical technique allows retrieving soliton content of nonlinear waves generated from arbitrary inputs in optical fibers [61], [62], [63], [64]. Thus, we reported input-power-dependent transitions from a quasi-linear regime, governed primarily by linear Talbot effect at low input powers, to the regimes of the nonlinear Talbot effect of soliton crystals and separated solitons (cf. Fig. 2, left). These regime transitions happened at well-defined threshold input powers, giving rise to the interpretation that the NLS changes the type of its solutions at these values.
@@
To the best of my knowledge, these were the first reported results to connect the temporal Talbot effect with the MI-related generation of soliton crystals and solitons in such detail. However, the results were derived rather heuristically from our SRBA observations. Many questions remained open, in particular, an explanation of the observed regimes and their transitions backed up by a concise theory of the nonlinear temporal Talbot, as the latter was not available at the time.
The main goal of this paper is to pave the way for the development of a unified and concise theory of the nonlinear temporal Talbot effect by integration of MI-related four-wave mixing (FWM) and cross-phase modulation (XPM) of frequency-comb lines, as well as their self-phase modulation (SPM), in the theory of the well-known linear temporal Talbot effect. This approach enables us to better understand various input-power-dependent regimes and regime transitions of the nonlinear temporal Talbot effect, including A-type breathers, soliton crystals, and soliton gas, opening new perspectives on the nonlinear-wave formation and their recurrence under FPUT.
@@
Thus, I build upon the results of Ref. [60] and study the transitions of the temporal Talbot effect and the physical mechanisms behind them using SRBA and bridging SRBA with a derived dispersion relation for frequency-comb lines. It allows me to reveal four input-power regimes: the regime of the linear temporal Talbot effect and the nonlinear regimes of A-type breathers, soliton crystals, and separated solitons (also called soliton gas due to a weak interaction between the solitons, [65]). I show that the (higher-order) interference of the discrete spatial modes of the frequency-comb lines originates from the temporal Talbot effect and governs the spatio-temporal distribution of the optical field in all regimes, allowing for the development of an exhaustive theory of the nonlinear temporal Talbot effect. The present study confirms the observation that frequency-comb lines can encode line-specific solitons, first reported in Ref. [60]. Additionally, I draw a difference between A-type breathers and soliton crystals. Thus, A-type breathers originate from regular FWM at quite low input powers, whereas soliton crystals stem from cascaded FWM at higher input powers. Exhibiting these characteristics, soliton crystals in single-pass optical fibers are no different from soliton crystals in cavities. Last but not least, I show that phase modulation depth is an important parameter that controls the type of nonlinear waves. Thus, a small modulation depth results in ABs under FPUT close to the Peregrine-soliton limit, whereas higher values of modulation depth lead to the formation of doubly periodic solutions such as A-type breathers and soliton crystals.
@@
To conclude, for decades, MI-induced nonlinear wave generation has been considered as a separate field of studies, giving rise to different types of NLS solutions. These solutions are important for the description and classification of various nonlinear waves in optical fibers and semiconductors. However, they provide only a limited understanding of the physics behind. My present study shows that a deeper and more sophisticated insight into the dynamics of a phase-modulated CW laser field stems from the perspective of the nonlinear temporal Talbot effect that integrates SPM and MI-related effects of FWM and XPM of frequency-comb lines in the theory of the well-known linear temporal Talbot effect. With that, it contributes to the development of an exhaustive theory of the nonlinear temporal Talbot effect and, thus, advances the fields of Nonlinear Optics and Wave Theory.
@@
The study is structured as follows. In Sec. 2, I present the system under study, a phase-modulated CW laser field in a single-mode optical fiber, and show that the operation happens beyond the standard AB theory. In Sec.3, I discuss the methodology for theoretical and numerical studies, including the SRBA. In the Results and Discussion section (Sec. 4), I present a dispersion relation for optical-frequency comb lines that includes group-velocity dispersion (GVD) and, with it, the footprint of the linear temporal Talbot effect, as well as SPM, XPM, and FWM, (Sec. 4.1); then, I bridge this dispersion relation with SRBA (Sec. 4.2); I discuss the transitions of the temporal Talbot effect from the linear to nonlinear regimes and draw the difference between the A-type breathers and soliton crystals (Sec. 4.3); I provide a confirmation that frequency-comb lines can encode line-specific solitons ([60], Sec. 4.4); and, finally, I discuss the impact of the modulation depth on the nonlinear-wave evolution (Sec. 4.5). I conclude and present a summary of the achieved results in Sec. 5.
@@ 2.1 The System Under Study: A Phase-Modulated Continuous-Wave Optical Field
The optical system I study here is depicted in Fig. 1 (right). It consists of a CW laser with emission at a central wavelength of $\lambda_{0}=1554.6~\text{nm}.$ The laser output is phase-modulated by the phase modulator (PM) with modulation frequency $\Omega=15.625~\text{GHz}$ and modulation depth $m.$ The phase-modulated CW field is propagated through a single-mode fiber (SMF) and recorded by an optical sampling oscilloscope (OSA) and optical spectrum analyzer (OSA). The nonlinear coefficient of the SMF is $\gamma=1.2~\text{(W}\cdot\text{km)}^{-1}$ and the GVD parameter $\beta_{2}=-23~\text{ps}^{2}/\text{km}.$ Please note that the choice of the of the fiber parameters and the value of the modulation frequency $\Omega=15.625~\text{GHz}$ is arbitrary. In particular, $\beta_{2}$ and $\Omega$ rescale the Talbot period $L_{T}$ [60], but do not alter the existence or nature of the regimes I will discuss below.
@@
Figure 2: Left: Soliton Radiation Beat Analysis of input-power dependent regimes in dB (cf. [60]). Green arrows mark ridges of spatio-temporally separated solitons. The yellow arrow marks the fundamental Talbot frequency $Z_{T}$ (Eq. 13). Right: Corresponding optical peak power $P(z)=|A(z,t=0)|^{2}$ in W. The PM modulation depth is $\text{m}=1.$ Dashed lines mark input-power dependent transitions of the nonlinear temporal Talbot effect.
According to Ref. [60], these regimes arise due to the temporal Talbot effect, a self-imaging effect in dispersive media such as optical fibers or semiconductors (cf. [26], [27]). The condition for this effect is satisfied by the initial frequency comb that is provided through the phase modulation of a CW laser field [19], [20], [21] (Fig. 1, right). The fiber dispersion acts as a unitary phase filter, inducing the break-up of the phase-modulated CW field into a train of identical pulses [23]. These pulses are subject to the Kerr nonlinearity of the fiber glass, which leads to the formation of the nonlinear regimes of soliton crystals and separated solitons. In Sects. 4.1, 4.2, 4.3, I provide a detailed explanation of how these regimes arise, building a bridge between the linear temporal Talbot effect and the nonlinear SPM and MI-driven FWM and XPM of frequency-comb lines. It allows for the development of a concise theory of the nonlinear temporal Talbot effect.
@@ 2.2 Operation Beyond the Established Akhmediev Breather Theory
Fig. 4 shows the MI gain $g(\Omega)$ and the corresponding AB parameters $a$ and $b$ for input powers studied in Ref. [60] and in more detail below. The AB parameters are fundamental to the AB solution of the NLS and are defined as $2a=(1-(\Omega/\Omega_{c})^{2})$ and $b=\sqrt{8a(1-2a)}$ with $\Omega_{c}=4\gamma P_{0}/|\beta_{2}|$ such that the maximum gain happens for $b=1$ and $a=1/4$ [8]. For $\beta_{2}=-23~\text{ps}^{2}/\text{km}$ used here, the AB parameters are $\leq 0$ and, thus, not satisfactory for an AB solution for the most significant input power values of $P_{0}=0.046~\text{W},$ $0.15~\text{W},$ and $0.27~\text{W}$ (cf. Fig. 3, top). Also, the higher-order MI does not come into question as a driving mechanism for our observations, as it requires a value of $a\geq 0.375$ [40] which is achievable only for low absolute values of GVD. Apart from that, the PM modulation frequency RF of $\Omega=15.625~\text{GHz}$ generates only a little MI gain far beyond its input-power dependent maxima. From these observations, I conclude that the results reported in Ref. [60] and studied here in more detail do not originate from ABs under FPUT recurrence despite their visual similarity ([3]). Rather, they are the result of the nonlinear temporal Talbot effect that incorporates SPM and MI-driven FWM and XPM of frequency-comb lines as discussed in Sec. 4.
@@
A more generalized class of NLS solutions constitutes the spatio-temporally periodic A-type and B-type breathers ([51], [52], [53]). In Ref. [54], the authors discussed that A-type breathers exist beyond the standard MI bandwidth of modulation frequencies, i.e., for higher modulation frequencies than the cut-off frequency of $g(\Omega)$ (Fig. 4, left), which gives rise to ”extraordinary” MI. Below, I will discuss a regime of the nonlinear temporal Talbot effect for input powers of $0.046~\text{W}<P_{0}\leq 0.15~\text{W}$ that hosts doubly-periodic waves with properties of A-type breathers (Sec. 4.3). Apparently, the PM modulation frequency does not need to exceed the cut-off frequency of $g(\Omega)$. It is also possible to generate A-type breathers within the MI bandwidth for a small value of $g(\Omega).$
@@
Soliton crystals, spatio-temporally periodic soliton compounds for $0.15~\text{W}<P_{0}\leq 0.27~\text{W},$ constitute a separate class of NLS solutions. The difference to A-type breathers lies in the underlying physics. As discussed in Secs. 4.1, 4.2, 4.5, the A-type breathers are generated by MI-driven regular FWM at quite low input powers and small PM modulation depths; soliton crystals originate from cascaded FWM at higher input powers and strong PM modulation depths. For higher input powers, $P>0.27~\text{W},$ soliton crystals dissolve into spatio-temporally separated solitons, i.e., soliton gas (cf. [65]).
@@
I will show that higher-order interference between the spatial modes of the frequency-comb lines driven by the temporal Talbot effect is the underlying mechanism for any type of the discussed NLS solutions (Sec. 4.3). Therefore, I conclude that an exhaustive consideration of nonlinear-wave evolution in optical fibers should be conducted from the perspective of the nonlinear temporal Talbot effect, rather than considering the linear temporal Talbot effect and SPM, as well as MI-driven effects of FWM and XPM separately, as usually done. Now, let us dive into more.